OpenStudy (ttop0816):

please help! Given the system of equations, what is the x-coordinate of the solution? y = 3 - 1/2x 3x + 4y = 1 22 -17/2 -11

OpenStudy (marsxtc):

-11. You have to isolate the y on the second equation. So it would be 3x+4y =1 3x-1 = -4y (3x-1)/-4 = y. Then you can set the equations to each other (3 - 1/2x)= (3x-1)/-4. Try doing that by yourself